OpenStudy (anonymous):

lim x->0 [x*cot((pi*x)/3)]. Why is the answer 3/pi ?

OpenStudy (lalaly):

do u know lhopitals rule?

OpenStudy (anonymous):

\[\frac{x\cot(\pi x)}{3}=\frac{x\cos(\pi x)}{\sin(\pi x)}\] \[=\frac{\pi x}{\pi \sin(\pi x)}\times \cos(\pi x)\] will work if you cannot use l'hopital

OpenStudy (lalaly):

lol i love lhopitals ... Nice @satellite73

OpenStudy (anonymous):

how bout them zeta functions?

OpenStudy (anonymous):

i misses a 3 somewhere in my last expression \[=\frac{\pi x}{3\pi \sin(\pi x)}\times \cos(\pi x)\] now let \(x\to 0\) and you should get \(\frac{\pi}{3}\) because \(\cos(0)=1\) and \(\lim_{x\to 0}\frac{\sin(x)}{x}=1\)

OpenStudy (anonymous):

damn i am making mistake after mistake

OpenStudy (anonymous):

i think the answer is \(\frac{1}{3\pi}\)

OpenStudy (anonymous):

its x*cot((pi*x)/3) not what you said, which was x*cot((pi*x))/3

OpenStudy (anonymous):

lets start here \[\frac{\pi x}{3\pi \sin(\pi x)}\times \cos(\pi x)\]

OpenStudy (anonymous):

\[\lim_{x\to 0}x\cot(\frac{\pi x}{3})\]

OpenStudy (anonymous):

your last one is correct

OpenStudy (anonymous):

ok same idea as last one

OpenStudy (anonymous):

\[x\cos(\frac{\pi x}{3})=\frac{x\cos(\frac{\pi x}{3})}{\sin(\frac{\pi x}{3})}\]

OpenStudy (anonymous):

multiply top and bottom by \(\frac{\pi}{3}\)

OpenStudy (anonymous):

\[\frac{\frac{\pi}{3}x\cos(\frac{\pi x}{3})}{\frac{\pi}{3}\sin(\frac{\pi x}{3})}\]

OpenStudy (anonymous):

then \[\lim_{x\to 0}\cos(\frac{\pi x}{3})=\cos(0)=1\]

OpenStudy (anonymous):

and \[\lim_{x\to 0}\frac{\frac{\pi}{3}x}{\sin(\frac{\pi x}{3})}=1\]

OpenStudy (anonymous):

leaving you with \[\frac{1}{\frac{\pi}{3}}=\frac{3}{\pi}\]

OpenStudy (anonymous):

why is this true?: xcos(πx/3)=xcos(πx/3)sin(πx/3)

OpenStudy (anonymous):

it isn't. you had cotangent, not cosine right?